2006 CNA Summer School

Probabilistic and Analytical Perspectives on Contemporary PDEs Center for Nonlinear Analysis Carnegie Mellon University

Pittsburgh, PA May 29–June 6, 2006

List of Participants

Vilen Abramov  Kent State University  vabramov@math.kent.edu
Tom Alberts  New York University  alberts@cims.nyu.edu
Robert Berry  University of Pittsburgh  rdb6@pitt.edu
David Bourne  University of Maryland  dbourne@math.umd.edu
Hsun-Hsien Chang  Carnegie Mellon University  hsunhsien@cmu.edu
Jung-Pin Chen  Purdue University  chen201@purdue.edu
Yanlai Chen  University of Minnesota  chen0731@umn.edu
Patrick Cheridito  Princeton  dito@princeton.edu
Huseyin Coskun  University of Iowa  hcoskun@math.uiowa.edu
Milica Cudina  Carnegie Mellon University  mcudina@gmail.com
Francois Delarue  University of Paris VII  delarue @math.jussieu.fr
Andras Domokos  California State University  domokos@csus.edu
Patrick Dondl  California Institute of Technology  pwd@caltech.edu
Bo Dong  University of Minnesota  bdong@math.umn.edu
Hongjie Dong  University of Chicago  hjdong@math.uchicago.edu
Federica Dragoni  Scuola Normale Superiore  f.dragoni@sns.it
Paula Cristina Faria  CMAF Lisbon University
and ESTG-IPLeiria  pfaria@estg.ipleiria.pt
Matthew Folz Simon Fraser University mfolz@sfu.ca
Edwin Bobga Fohtung  St. Petersburg State Polytechnic University fohtung ed@physics.org
Jasmine Foo  Brown University  jfoo@dam.brown.edu
Savannah Gomes  University of Toronto  savannah.gomes@utoronto.ca
Ilie Grigorescu  University of Miami  igrigore@math.miami.edu
Johnny Guzman  University of Minnesota  guzma033@umn.edu
Boubakari Ibrahimou  University of South Florida  bibrahim@mail.usf.edu
Prabhu Janakiraman  University of Illinois  pjanakir@math.uiuc.edu
Juhi Jang  Brown University  juhijang@math.brown.edu
Shaine Joseph  University of Missouri  shaine joseph@yahoo.com
Vladimir Kamotski  University of Bath  v.kamotski@bath.ac.uk
Young-Heon Kim  Fields Institute  ykim@fields.utoronto.ca
Siu Tang Leung  Princeton University  siutang@princeton.edu
Dmitriy Leykekhman  Rice University  dmitriy@caam.rice.edu
Hengguang Li  Penn State University  li h@math.psu.edu
Joan Lind  Cornell University  jlind@math.cornell.edu
Chin-Yueh Liu  University of Minnesota  liux0518@umn.edu
Tao Luo  Purdue University  tluo@math.purdue.edu
Sheng Lynn  Carnegie Mellon University  iangmiao@yahoo.com
Antonio Mastroberardino  SUNY, Buffalo  am59@buffalo.edu
Ivan Matic  University of California, Berkeley  matic@math.berkeley.edu
Stephane Menozzi  Universite Denis Diderot, Paris VII  menozzi@math.jussieu.fr
Akil Narayan  Brown University  anaray@dam.brown.edu
Robert Neel  Columbia University  neel@math.columbia.edu
Daniel Onofrei  Worcester Polytechnic Institute  onofrei @wpi.edu
Eduardo Osorio  Rutgers  eduardos@eden.rutgers.edu
Seshadev Padhi  Mississippi State University  sp377@ra.msstate.edu
Mahendra Panthee  Instituto Superior Tecnico, Portugal  mpanthee@math.ist.utl.pt
Jinhae Park  University of Minnesota  jinhae@math.umn.edu
Swabhimita Patel  University of Houston  spatel@math.uh.edu
Jonathan Pietarilla Graham University of Colorado
National Center for Atmospheric Research jpietarilagraham@mailaps.org
Remy Praz  Carnegie Mellon University/EPFL  remy.praz@epfl.ch
Alex Rand  Carnegie Mellon University  arand@cmu.edu
Reza Rastegar  Southern Illinois University  ras reza@yahoo.com
Marco Roque-Sol  Texas a&M University  roquesol@math.tamu.edu
Russell Schwab  University of Texas, Austin  rschwab@math.utexas.edu
Jaffar Ali Shahul Hameed Mississippi Sate University  j2415@ra.msstate.edu
Ravi Srinivasan  Brown University  rav@dam.brown.edu
Fang Wan  Purdue University  fwan@purdue.edu
Fang Wan  Purdue University  fwan@purdue.edu
Hao Wang  University of Oregon  haowang@uoregon.edu
David White  Cornell University  dwhite@math.cornell.edu
Robert Wooster  University of Connecticut  wooster@math.uconn.edu
Dongshen Wu  Michigan State University  wudongsh@ms.edu
Xiaoqin Wu  Mississippi State University  xw2@msstate.edu
Jianhua Xu  Carnegie Mellon University  jianhuax@andrew.cmu.edu
Guangri Xue  Penn State University  xue@math.psu.edu
KiHyun Yun  Carnegie Mellon University  kyun@andrew.cmu.edu
Anne Yust  Carnegie Mellon University  aey@andrew.cmu.edu
Chensong Zhang  University of Maryland  zhangcs@math.umd.edu
Hang Zhang  University of Minnesota  hzhang@math.umn.edu
Jian Zhang  Penn State University  zhang j@math.psu.edu
Lei Zhang  California Institute of Technology  zhanglei@acm.caltech.edu
Yihui Zhou  University of Maryland  yihui1@umbc.edu
Yun Zhou  Columbia University  zhou@math.columbia.edu
Gordan Zitkovic  University of Texas at Austin  gordanz@math.utexas.edu

2